# NestQuant: Nested Lattice Quantization for Matrix Products and LLMs

### v1

@@ Abstract
Post-training quantization (PTQ) has emerged as a critical technique for efficient deployment of large language models (LLMs). This work proposes NestQuant, a novel PTQ scheme for weights and activations that is based on self-similar nested lattices. Recent work have mathematically shown such quantizers to be information-theoretically optimal for low-precision matrix multiplication. We implement a practical low-complexity version of NestQuant based on Gosset lattice, making it a drop-in quantizer for any matrix multiplication step (e.g., in self-attention, MLP etc). For example, NestQuant quantizes weights, KV-cache, and activations of Llama-3-8B to 4 bits, achieving perplexity of 6.6 on wikitext2. This represents more than 55% reduction in perplexity gap with respect to unquantized model (perplexity of 6.14) compared to state-of-the-art Meta’s SpinQuant (perplexity 7.3). Comparisons on various LLM evaluation benchmarks also show a reduction in performance degradation induced by quantization.
@@ 1 Introduction
Figure 1: Perplexity of quantized models for three regimes (weight-only, weights + KV cache, end-to-end) on wikitext2 vs number of bits per entry.
@@
The second goal of PTQ is to accelerate inference. In LLMs most of the compute is spent multiplying matrices. Multiplying a pair of such matrices requires $2n^{3}$ FLOPs and ${3\over 8}Rn^{2}$ bytes to exchange between the core and memory (here and below $R$ designates the number of bits required to store each entry of a vector/matrix). So when matrices are large (such as during the pre-fill phase when the prompt is processed) the GPU is compute-bound, while when $n$ is small (such as during generation) the GPU becomes memory-bound. To achieve this goal one needs to reduce $R$ by quantizing both weights and the KV cache.
@@
While quantization of weights to $R=3,4$ and even $R=2$ bits has been achieved with minimal loss of quality, quantization of KV cache and activations has been much more challenging. Popular algorithms for full quantization are LLM.int8(), SmoothQuant and SpinQuant (see references in Section 2.2), the latter having state-of-the-art performance. This work proposes an alternative algorithm (NestQuant) for quantizing weights, KV-cache, and activations. The algorithm is motivated by recent theoretical work on approximate matrix multiplication and follows several classical ideas such as the Conway-Sloane algorithm for the Gosset lattice.
@@ 1.1 Summary of results
The NestQuant algorithm is described in Section 4. NestQuant is a generic drop-in replacement for any matrix multiplication. Its performance for synthetic random Gaussian matrices comes pretty close to information-theoretic limits (see Fig. 3) and significantly outperforms uniform quantization employed by SpinQuant. Switching from a scalar (uniform) quantization to vector quantization requires some price to pay computationally (Section C), however, among vector quantizers NestQuant is rather economical as it operates in dimension 8 and leverages an elegant algorithm of [1].
Applying NestQuant to quantizing an actual LLM (Llama-3-8B) shows massive end-to-end improvement: Fig. 1 shows a significant reduction of perplexity compared to SpinQuant; and Table 1 confirms enhanced performance on standard LLM benchmarks.
@@
Thus, we believe that NestQuant offers an excellent alternative to other algorithms. It quantizes weights, KV-cache and activations, achieves significant improvement on both synthetic and real data.
@@ 2.1 Information-theoretic quantization
Rate $R$ quantization of an information source $X$ in $\mathbb{R}^{n}$ is the operation of encoding it to $nR$ bits, from which a decoder can produce a reconstruction $\hat{X}\in\mathbb{R}^{n}$ that has small distortion with respect to $X$. The most popular distortion criterion is the quadratic loss, where the expected distortion is defined as $D=\frac{1}{n}\mathbb{E}\|X-\hat{X}\|^{2}$, and here we restrict attention to this loss. Characterization of the optimal tradeoff between $R$ and $D$ is a classic topic in information theory, e.g. [2, Part V].
@@
One way for constructing these universal quantizers is based on lattices [3] that admit much more structure than more classical random codes (and $\epsilon$-nets).
Arguably, the most notable lattice-based quantization scheme is the family of Voronoi codes [1], which we use in this work.
How does one convert a quantizer adapted to Gaussian inputs to work (with the same guaranteed loss) on non-Gaussian data? In a nutshell, the idea is simple: if $U$ is chosen to be a random $n\times n$ orthogonal matrix then the entries of $UX$ will be distributed as iid Gaussian [4]. This idea of applying random rotations to smooth out the distribution of the quantizer’s input may be viewed as a special case of high-dimensional companding [5], and has been applied for image compression [6], and as a potential replacement for dithering [7], to name a few.
In the context of LLMs, the goal in quantization is slightly different since quantization is used to facilitate approximate matrix multiplication with reduced burden on the memory bandwidth. For example, when quantizing two vectors $X,Y\in\mathbb{R}^{n}$ the goal is not to approximate them but to approximate their inner product. Recently, information-theoretic characterization of this task was completed in [8]. Specifically, the authors show that if $X,Y\sim\mathcal{N}(0,I_{n})$ are independent then for any algorithm operating on the basis of rate-$R$ quantized representations of $X$ and $Y$ we must have
@@
where
$\Gamma(R)=\begin{cases}1-\left(1-(2\cdot 2^{-2R^{*}}-2^{-4R^{*}})\right)\frac{R}{R^{*}}&R<R^{*}\\ 2\cdot 2^{-2R}-2^{-4R}&R\geq R^{*}\end{cases}.$ (2)
and $R^{*}\approx 0.906$ is a solution to a certain transcendental fixed-point equation.
@@
In this work we show that with appropriate tweaks Voronoi codes indeed can result in practical fast and efficient algorithms for LLM quantization. We emphasize that most of our work is on simply developing a drop-in replacement for quantized matrix product and as such is not specific to LLMs.
The idea of applying random rotations to “Gaussianize” inputs in the context of approximate inner product computation is even more natural than in the standard quantization. Indeed, since one is only interested in the inner product, not vectors themselves, one does not need to store (even a seed used to generate the) random orthogonal matrix. This has been long exploited in locality-sensitive hashing (LSH) algorithms,***In LSH, one typically performs several random projections of the vector and quantizes them. This is equivalent to performing random rotation and quantizing only a small number of entries of the rotated vector. which can be viewed as an (extremely low rate) quantizers for inner product approximation [9, 10, 11]. Unsurprisingly, as we will see next, random rotations have also been found quite useful for quantizing LLMs.
@@ 2.2 LLM quantization
One of the directions of prior research on LLM quantization is addressing the issue of activation outliers that hinder the quantization quality. These outliers are present in certain dimensions of activations, weights and KV cache. In LLM.int8() of [12], these outlier dimension are kept unquantized. In SmoothQuant [13] authors balance the scale of outliers between weights and activations by modifying LayerNorm’s diagonal matrices.
Going to random rotations, by rewriting matrix product $AB=(AU)(U^{\top}B)$ for an orthogonal matrix $U$, one gets matrices with much more Gaussian entries (few outliers) and can apply standard quantization algorithms. Some of the multiplications by $U$ can be merged with the weights (i.e. do not require additional runtime FLOPs), while the rest are applied at runtime. For the latter, matrices $U$ should have structure to enable fast multiplication. For example, QuaRot [14] uses randomized Hadamard matrices as coordinate transformations, which can be applied to a vector of size $n$ in $O(n\log n)$ additions. SpinQuant [15] uses a rotation parametrization with four orthogonal matrices $R_{1}$, $R_{2}$, $R_{3}$, $R_{4}$, where $R_{1}$ and $R_{2}$ can be arbitrary orthonormal matrices, and $R_{3}$ and $R_{4}$ should have a fast multiplication algorithm. The authors use Cayley SGD [16] to optimize $R_{1}$ and $R_{2}$ for minimization of the quantization error, while the matrices $R_{3}$ and $R_{4}$ are chosen to be random Hadamard.
Starting from LLM.int8() most of the schemes used uniform quantization (i.e. where a floating point vector simply rounded to a nearest integer after an appropriate rescaling). To the best of our knowledge, so far non-scalar quantization has only been used for weight-only compression for LLMs in QuIP# [17], which uses E8P codebook for 2-bit quantization, and applies Resdidual Vector Quantization [18] to get a 4-bit quantization scheme; and QTIP [19] which uses trelis codebook. Unfortunately these methods appear to be too expensive to apply then in runtime, perhaps explaining why non-uniform quantization for activations and KV-cache was not attempted before this work.
Finally, when quantizing weight matrices, one may notice that MSE distortion loss should be replaced by a weighted-MSE loss dependent on the statistics of the incoming activations. We refer to this type of algorithms as LDLQ, following authors of QuIP [20], QuIP# [17] and GPTQ [21]. Applying LDLQ-modified NestQuant does lead to improved performance, see results in Section 5.
@@ 3 Outline of NestQuant approach
When designing a quantizer, one needs to make some assumptions about the distribution of the source that will be fed to it. While weights (and sometimes activations) can be well-approximated by Gaussians, their magnitude are wildly varied. Thus, one employs two ideas: normalization and random rotation.
Normalization: In most of the literature, the normalization is done by taking an input vector of large dimension $n$ (e.g. $n=4096$ for Llama-3), dividing by the $L_{\infty}$ norm to get entries to be in $[-1,1]$ and then applying uniform quantization. This is suboptimal for two reasons: one is that uniform quantization induces error that is distributed uniformly on the small cube, which is suboptimal from the MSE point of view. Second reason, much more serious, is known as the shaping gain and demonstrated on Fig. 2. When entries of the vector are Gaussian, it will typically lie inside the black circle. Thus those grid elements outside of it will almost never be used, wasting bitspace.
@@
Random rotation: When input to the quantizer significantly differs from the presumed model (of iid Gaussians), performance can become quite poor. As discussed previously, multiplying by a random orthoghonal matrix $U$ provides an excellent fix. Specifically, $UX$ vector becomes uniform on the $n$-sphere of radius $\sqrt{n}$, and small chunks of this vector have distribution very close to Gaussian iid. In particular, the total variation between any subset of $d$ coordinates and $\mathcal{N}(0,I_{d})$ is $O(d^{2}/n)$ [4], such that for $d=o(\sqrt{n})$ what we quantize is effectively iid Gaussian.
Complexity of lattice quantization: In order to explain our choice of nested lattice quantizer, we need to carefully balance several requirements. One of the key ones is complexity. It is known that finding (even approximating) a nearest lattice point is a well-known cryptographic assumption [22]. Thus, we are not suggesting to directly operate on $n$-dimensional lattices. Instead, we partition the $n$-vector into sections, each of dimension $d$ and apply lattice quantization to $d$-subvectors. Equivalently, our vector quantizers for $\mathbb{R}^{n}$ are constructed as Cartesian products of vector quantizers of small dimension $d$ (we will take $d=8$ for all experiments).
@@
where $\mathcal{V}_{\Lambda}\subset\mathbb{R}^{d}$ is the Voronoi region of $\Lambda$, consisting of all point in $\mathbb{R}^{d}$ closer to 00 than to any other lattice point in $\Lambda$. This quantity corresponds to the MSE when $\Lambda$ is normalized to have unit covolume and is then used as a quantizer. It is well known that the optimal (smallest) NSM among all lattices in $\mathbb{R}^{d}$ approaches $\frac{1}{2\pi e}$ from above as $d$ increases [3]. Furthermore, for $d=1$ we get $G(\mathbb{Z})=\frac{1}{12}$. Consequently, in terms of granular error, by using high-dimensional lattices instead of the simple scalar quantizer based on $\mathbb{Z}$ we can already gain a factor of $\frac{2\pi e}{12}\approx 1.42329$ in performance (the Gosset lattice achieves $1.22$ gain).
Notice, however, that representing $x$ as $Q_{\Lambda}(x)$, the nearest lattice point in $\Lambda$, requires infinitely many bits, since the lattice is infinite. Since we only have $2^{dR}$ bitstrings to allocate, we need to select a subset of $\Lambda$ that will be actually used. Selection of $\mathcal{S}$ so that $|\Lambda\cap\mathcal{S}|=2^{dR}$ is called shaping. If $Q_{\Lambda}(x)\in\mathcal{S}$, then the quantization error $x-Q_{\Lambda}(x)$ will be in $\mathcal{V}_{\Lambda}$ and we will only suffer from a granular error. However, when $Q_{\Lambda}(x)\notin\mathcal{S}$ the quantization error is no longer in $\mathcal{V}_{\Lambda}$ and may be have much greater magnitude than a typical granular error. We refer to those type of errors, where $Q_{\Lambda}(x)\notin\mathcal{S}$, as overload errors.
@@
Denote by $\mu=\mathcal{N}(0,I_{d})$ the standard Gaussian measure. Since we need $2^{dR}=|\Lambda\cap\mathcal{S}|\approx\frac{\mathrm{vol}(\mathcal{S})}{\mathrm{covol}(\Lambda)}$, a good shaping region $\mathcal{S}$ maximizes $\mu(\mathcal{S})$, which in turn minimizes the overload probability that is approximated by $1-\mu(\mathcal{S})$, under a volume constraint. Clearly, the optimal $\mathcal{S}$ under this criterion is $r\mathcal{B}$ where =$\mathcal{B}=\{x\in\mathbb{R}^{d}~{}:~{}\|x\|\leq r_{\text{eff}}(1)\}$ is a Euclidean ball with radius $r_{\text{eff}}(1)$ chosen such that $\mathrm{vol}(\mathcal{B})=1$, and $r$ is chosen such that $\mathrm{vol}(\mathcal{S})=r^{d}$ satisfies the required volume constraint. Unfortunately, for $d>1$ the codebook $\mathcal{C}=\Lambda\cap r\mathcal{B}$ loses much of the lattice structure, and does not admit an efficient enumeration, and consequently encoding and decoding require using a lookup table (LUT). QuIP# used this approach with $\Lambda=E_{8}$ (same as we do) and $\mathcal{S}=r\mathcal{B}$. However, this seems to only be possible for quantizing weights and not activations as complexity makes runtime implementation too slow.We note that QuIP# cleverly exploits symmetries of $E_{8}$ to show that an $R=2$ bit quantizer can be implemented using an LUT of size $2^{d\frac{R}{2}}=2^{8}$, but we believe this is still too slow, and furthermore does not naturally extend to different quantization rates.
@@
Voronoi codes/nested lattice codes: In Voronoi codes [1] the same lattice $\Lambda$ is used for both quantization and shaping. In particular, the shaping region is taken as $\mathcal{S}=2^{R}\mathcal{V}_{\Lambda}$, where $2^{R}$ is an integer. As elaborated below, if $Q_{\Lambda}(x)$ admits an efficient implementation, one can efficiently perform encoding and decoding to the codebook $\mathcal{C}=\Lambda\cap(2^{R}\mathcal{V}_{\Lambda})\cong\Lambda/2^{R}\Lambda$. Moreover, in stark contrast to ball-based shaping, the encoding and decoding complexity does not depend on $R$.
@@
In this work, we use the Gosset lattice ($E_{8}$) that satisfies all these properties. It has a fast decoding algorithm (Algorithm 5), its NSM is $\approx 0.0716821\approx 1.2243\frac{1}{2\pi e}$ [23], and its Gaussian mass $\mu(r\mathcal{V}_{E_{8}})$ is very close to $\mu(r\mathcal{B})$ (the Gosset lattice has unit covolume, so $\mathrm{vol}(r\mathcal{V}_{E_{8}})=\mathrm{vol}(r\mathcal{B})$). The last point is illustrated in Figure 6, where the large loss for cubic shaping with respect to lattice shaping is also evident. The gap between Voronoi/Ball shaping and cubic shaping becomes much more significant as the dimension increases. This follows since for large $d$ we have that for $X\sim\mu=\mu_{d}$ the $\ell_{\infty}$ norm $\|X\|_{\infty}$ concentrates around $\sqrt{2\ln{d}}$. Thus, for $r<2\sqrt{2\ln{d}}$ we have that $\mu(r\mathrm{CUBE})\to 0$, whereas for any $r>\frac{\sqrt{d}}{r_{\text{eff}}(1)}=\sqrt{2\pi e}(1+o(1))$ we have that $\mu(r\mathrm{B})\to 1$. Note that SpinQuant uses high-dimensional cubic shaping, and therefore its MSE distortion suffers a $O(\ln d)$ multiplicative gap with respect to the optimal distortion.

### v3

@@ Abstract
Post-training quantization (PTQ) has emerged as a critical technique for efficient deployment of large language models (LLMs). This work proposes NestQuant, a novel PTQ scheme for weights and activations that is based on self-similar nested lattices. Recent works have mathematically shown such quantizers to be information-theoretically optimal for low-precision matrix multiplication. We implement a practical low-complexity version of NestQuant based on Gosset lattice, making it a drop-in quantizer for any matrix multiplication step (e.g., in self-attention, MLP etc). For example, NestQuant quantizes weights, KV-cache, and activations of Llama-3-8B to 4 bits, achieving perplexity of 6.6 on wikitext2. This represents more than 55% reduction in perplexity gap with respect to unquantized model (perplexity of 6.14) compared to state-of-the-art Meta’s SpinQuant (perplexity 7.3), OstQuant (7.3) and QuaRot (8.2). Comparisons on bigger models (up to 70B) and on various LLM evaluation benchmarks confirm uniform superiority of NestQuant.
@@ 1 Introduction
Figure 1: Perplexity vs number of bits/entry in quantized Llama-3-8B for three regimes (weight-only, weights + KV cache, end-to-end) on wikitext2 at 2048 context.
@@
The second goal of PTQ is to accelerate inference. In LLMs most of the compute is spent multiplying matrices. Multiplying a pair of $n\times n$ matrices requires $2n^{3}$ FLOPs and ${\frac{3}{8}}Rn^{2}$ bytes to exchange between the core and memory (here and below $R$ designates the number of bits required to store each entry of a vector/matrix). So when matrices are large (such as during the pre-fill phase when the prompt is processed) the GPU is compute-bound, while when $n$ is small (such as during generation) the GPU becomes memory-bound. One therefore needs to reduce $R$ by quantizing both weights and the KV cache, thereby alleviating the memory bandwidth bottleneck during generation.
@@
While quantization of weights to $R=3,4$ and even $R=2$ bits has been achieved with minimal loss of quality, quantization of KV cache and activations has been much more challenging. Popular algorithms for full quantization are LLM.int8(), SmoothQuant and SpinQuant (see references in Section 2.2), the latter having state-of-the-art performance. This work proposes an alternative algorithm (NestQuant) for quantizing weights, KV-cache, and activations. The algorithm is motivated by recent theoretical work on approximate matrix multiplication and follows several classical ideas such as Conway and Sloane’s Voronoi codes and their algorithms for finding the closest lattice vector in some canonical lattices in low/moderate dimensions.
Throughout this paper, NestQuant is described with the 888-dimensional Gosset lattice, but the framework is general and can be similarly implemented using any other base lattice, in any dimension, provided that a low-complexity closest lattice vector algorithm exists for this lattice.
@@ 1.1 Summary of results
The NestQuant algorithm is described in Section 4. NestQuant is a generic drop-in replacement for any matrix multiplication. Its performance for synthetic random Gaussian matrices comes pretty close to information-theoretic limits (see Fig. 3) and significantly outperforms uniform quantization employed by SpinQuant. Switching from a scalar (uniform) quantization to vector quantization requires some price to pay computationally (Section E.1), however, among vector quantizers NestQuant is rather economical as it builds upon the Voronoi Codes framework of (ConwaySloane83, ). We mention that in the presence of activation quantization, it is important to quantize weights properly – an innovation we call QA-LDLQ (Section 4.5).
Applying NestQuant to quantizing an actual LLMs (Llama 2 and 3 with parameter count ranging from 1B to 70B) shows massive end-to-end improvement: Fig. 1 shows a significant reduction of perplexity compared to the previous methods; Table 1 confirm enhanced performance on standard LLM benchmarks; and Table 2 show NestQuant stays consistently on top for different model sizes. In fact, NestQuant in full quantization (weights, activations, and KV cache) outperforms SOTA results with quantized weights and activations (but not KV cache).
@@
Thus, we believe that NestQuant offers an excellent alternative to other algorithms. It quantizes weights, KV-cache and activations, and achieves significant improvement on both synthetic and real data.
@@ 2.1 Information-theoretic quantization
Rate $R$ quantization of an information source $X$ in $\mathbb{R}^{n}$ is the operation of encoding it to $nR$ bits, from which a decoder can produce a reconstruction $\hat{X}\in\mathbb{R}^{n}$ that has small distortion with respect to $X$. The most popular distortion criterion is the quadratic loss, where the expected distortion is defined as $D=\frac{1}{n}\mathbb{E}\|X-\hat{X}\|^{2}$, and here we restrict attention to this loss. Characterization of the optimal tradeoff between $R$ and $D$ is a classic topic in information theory, e.g. (PWbook24, , Part V).
@@
One way for constructing these universal quantizers is based on lattices ramiBook that admit much more structure than more classical random codes (and $\epsilon$-nets).
Arguably, the most notable lattice-based quantization scheme is the family of Voronoi codes ConwaySloane83 , which we use in this work.
How does one convert a quantizer adapted to Gaussian inputs to work (with the same guaranteed loss) on non-Gaussian data? In a nutshell, the idea is simple: if $U$ is chosen to be a random $n\times n$ orthogonal matrix then the entries of $UX$ will be distributed as iid Gaussian stam1982limit . This idea of applying random rotations to smooth out the distribution of the quantizer’s input may be viewed as a special case of high-dimensional companding gersho1979asymptotically , and has been been used in various applications, such as image compression hung1998multidimensional , and as a potential replacement for dithering hadad2016dithered , to name a few.
[…]
@@
where $\widehat{X^{\top}Y}$ is the reconstructed inner product, and
$\Gamma(R)=\begin{cases}1-\left(1-(2\cdot 2^{-2R^{*}}-2^{-4R^{*}})\right)\frac{R}{R^{*}}&R<R^{*}\\ 2\cdot 2^{-2R}-2^{-4R}&R\geq R^{*}\end{cases},$ (2)
where $R^{*}\approx 0.906$ is a solution to a certain transcendental fixed-point equation.
@@
Remark 2.1.
Using optimal quantizers, in terms of achieving $D(R)$, for quantizing $X$ and $Y$, and then computing the inner product of the quantized vectors (maybe with some MMSE scaling) will not necessarily lead to the optimal performance $\Gamma(R)$. The reason is that the covariance matrix of each vector’s quantization error affects the inner product distortion through its trace, and its Frobenius norm op2024 . For optimal lattice quantizers in any dimension, the Frobenius norm is guaranteed to be minimal zamir1996lattice , and consequently, using lattice quantizers seems essential for achieving $\Gamma(R)$.
In this work we show that with appropriate tweaks Voronoi codes can indeed result in practical fast and efficient algorithms for LLM quantization. We emphasize that most of our work is on simply developing a drop-in replacement for quantized matrix product and as such is not specific to LLMs.
The idea of applying random rotations to “Gaussianize” inputs in the context of approximate inner product computation is even more natural than in the standard quantization. Indeed, since one is only interested in the inner product, not vectors themselves, one does not need to store (even a seed used to generate the) random orthogonal matrix. This has been long exploited in locality-sensitive hashing (LSH) algorithms,***In LSH, one typically performs several random projections of the vector and quantizes them. This is equivalent to performing random rotation and quantizing only a small number of entries of the rotated vector. which can be viewed as an (extremely low rate) quantizers for inner product approximation charikar2002similarity ; datar2004locality ; andoni2008near . Unsurprisingly, as we will see next, random rotations have also been found quite useful for quantizing LLMs.
@@ 2.2 LLM quantization
One of the directions of prior research on LLM quantization is addressing the issue of activation outliers that hinder the quantization quality. These outliers are present in certain dimensions of activations, weights and KV cache. In LLM.int8() of dettmers2022 , these outlier dimension are kept unquantized. In SmoothQuant xiao2024 authors balance the scale of outliers between weights and activations by modifying LayerNorm’s diagonal matrices.
Going to random rotations, by rewriting matrix product $AB=(AU)(U^{\top}B)$ for an orthogonal matrix $U$, one gets matrices with much more Gaussian entries (few outliers) and can apply standard quantization algorithms. Some of the multiplications by $U$ can be merged with the weights (i.e. do not require additional runtime FLOPs), while the rest are applied at runtime. For the latter, matrices $U$ should have structure to enable fast multiplication. For example, QuaRot ashkboos2024 uses randomized Hadamard matrices as coordinate transformations, which can be applied to a vector of size $n$ in $O(n\log n)$ additions. SpinQuant liu2024 uses a rotation parametrization with four orthogonal matrices $R_{1}$, $R_{2}$, $R_{3}$, $R_{4}$, where $R_{1}$ and $R_{2}$ can be arbitrary orthonormal matrices, and $R_{3}$ and $R_{4}$ should have a fast multiplication algorithm. The authors use Cayley SGD li2020 to optimize $R_{1}$ and $R_{2}$ for minimization of the quantization error, while the matrices $R_{3}$ and $R_{4}$ are chosen to be random Hadamard.
Starting from LLM.int8() most of the schemes used uniform quantization (i.e. where a floating point vector simply rounded to a nearest integer after an appropriate rescaling). To the best of our knowledge, so far non-scalar quantization has only been used for weight-only compression for LLMs in QuIP# tseng2024 , which uses E8P codebook for 2-bit quantization, and applies Resdidual Vector Quantization Juang1982MultipleSV to get a 4-bit quantization scheme; and QTIP tseng2024qtip which uses trelis codebook. Unfortunately these methods appear to be too expensive to apply in runtime, perhaps explaining why non-uniform quantization for activations and KV-cache was not attempted before this work.
Finally, when quantizing weight matrices, one may notice that MSE distortion loss should be replaced by a weighted-MSE loss dependent on the statistics of the incoming activations. We refer to this type of algorithms as LDLQ, following authors of QuIP (chee2024, ), QuIP# (tseng2024, ) and GPTQ (frantar2023, ). In the presence of activation quantization, however, the LDLQ needs to be modified (see Section 4.5 introducing QA-LDLQ). Equipping NestQuant with QA-LDLQ incurs significant improvement; see results in Section 5.
@@ 3 Outline of NestQuant approach
When designing a quantizer, one needs to make some assumptions about the distribution of the source that will be fed to it. While weights (and sometimes activations) can be well-approximated by Gaussians, their magnitudes are wildly varied. Thus, one employs two ideas: normalization and random rotation.
Normalization: In most of the literature, the normalization is done by taking an input vector of large dimension $n$ (e.g. $n=4096$ for Llama-3), dividing by the $L_{\infty}$ norm to get entries to be in [1,1][-1,1][ - 1 , 1 ] and then applying uniform quantization. This is suboptimal for two reasons: one is that uniform quantization induces error that is distributed uniformly on the small cube, which is suboptimal from the MSE point of view. Second reason, much more serious, is known as the shaping gain and demonstrated on Fig. 2. When entries of the vector are Gaussian, it will typically lie inside the black circle. Thus those grid elements outside of it will almost never be used, wasting bitspace.
@@
Random rotation: When input to the quantizer significantly differs from the presumed model (of iid Gaussians), performance can become quite poor. As discussed previously, multiplying by a random orthoghonal matrix $U$ provides an excellent fix. Specifically, $UX$ vector becomes uniform on the $n$-sphere of radius $\sqrt{n}$, and small chunks of this vector have distribution very close to Gaussian iid. In particular, the total variation between any subset of $d$ coordinates and $\mathcal{N}(0,I_{d})$ is $O(d^{2}/n)$ stam1982limit , such that for $d=o(\sqrt{n})$ what we quantize is effectively iid Gaussian.
Lattices: A lattice $\Lambda\subset\mathbb{R}^{d}$ is a discrete subgroup of $\mathbb{R}^{d}$. Any lattice $\Lambda\subset\mathbb{R}^{d}$ has a (non-unique) generating matrix $G\in\mathbb{R}^{d\times d}$, such that $\Lambda=G\mathbb{Z}^{d}$ (that is, $\Lambda$ is the integer span of the columns of $G$). We define the nearest neighbor quantizer $Q_{\Lambda}:\mathbb{R}^{d}\to\Lambda$ as
$Q_{\Lambda}(x)=\operatorname*{\arg\!\min}_{\lambda\in\Lambda}\|x-\lambda\|,$ (3)
where ties are broken arbitrarily, but in a systematic manner. The Voronoi region $\mathcal{V}_{\Lambda}$ is defined as the set of all points in $\mathbb{R}^{d}$ that are closer to 0 than to any other lattice point
$\mathcal{V}_{\Lambda}=\left\{x\in\mathbb{R}^{d}~:~Q_{\Lambda}(x)=0\right\}.$ (4)
The covolume of a lattice is defined as $\mathrm{covol}(\Lambda)=|\det G|=\mathrm{vol}(\mathcal{V}_{\Lambda})$. We say that a lattice $\Lambda_{c}\subset\mathbb{R}^{d}$ is nested in the lattice $\Lambda_{f}\subset\mathbb{R}^{d}$, if $\Lambda_{c}\subset\Lambda_{f}$. Note that for any integer $q\geq 2$ we have that $q\Lambda\subset\Lambda$, and that $\Lambda/q\Lambda\cong(\mathbb{Z}/q\mathbb{Z})^{d}$. For an introduction to lattices and their role in quantization, see ramiBook .
Complexity of lattice quantization: In order to explain our choice of nested lattice quantizer, we need to carefully balance several requirements. One of the key ones is complexity. It is known that finding (even approximating) a nearest lattice point is a well-known cryptographic assumption DBLP:journals/combinatorica/DinurKRS03 . Thus, we are not suggesting to directly operate on $n$-dimensional lattices. Instead, we partition the $n$-vector into sections, each of dimension $d$ and apply lattice quantization to $d$-subvectors. Equivalently, our vector quantizers for $\mathbb{R}^{n}$ are constructed as Cartesian products of vector quantizers of small dimension $d$ (we will take $d=8$ for all experiments).
@@
This quantity corresponds to the MSE when $\Lambda$ is normalized to have unit covolume and is then used as a quantizer. It is well known that the optimal (smallest) NSM among all lattices in $\mathbb{R}^{d}$ approaches $\frac{1}{2\pi e}$ from above as $d$ increases ramiBook . Furthermore, for $d=1$ we get $G(\mathbb{Z})=\frac{1}{12}$. Consequently, in terms of granular error, by using high-dimensional lattices instead of the simple scalar quantizer based on $\mathbb{Z}$ we can already gain a factor of $\frac{2\pi e}{12}\approx 1.42329$ in performance (the Gosset lattice achieves 1.161.161.16 gain).
Notice, however, that representing $x$ as $Q_{\Lambda}(x)$, the nearest lattice point in $\Lambda$, requires infinitely many bits, since the lattice is infinite. Since we only have $2^{dR}$ bitstrings to allocate, we need to select a subset of $\Lambda$ that will be actually used. Selection of a body $\mathcal{S}\subset\mathbb{R}^{d}$ so that $|\Lambda\cap\mathcal{S}|=2^{dR}$ is called shaping. If $Q_{\Lambda}(x)\in\mathcal{S}$, then the quantization error $x-Q_{\Lambda}(x)$ will be in $\mathcal{V}_{\Lambda}$ and we will only suffer from a granular error. However, when $Q_{\Lambda}(x)\notin\mathcal{S}$ the quantization error is no longer in $\mathcal{V}_{\Lambda}$ and may be have much greater magnitude than a typical granular error. We refer to those type of errors, where $Q_{\Lambda}(x)\notin\mathcal{S}$, as overload errors.
@@
Denote by $\mu=\mathcal{N}(0,I_{d})$ the standard Gaussian measure. Since we need $2^{dR}=|\Lambda\cap\mathcal{S}|\approx\frac{\mathrm{vol}(\mathcal{S})}{\mathrm{covol}(\Lambda)}$, a good shaping region $\mathcal{S}$ maximizes $\mu(\mathcal{S})$, which in turn minimizes the overload probability that is approximated by $1-\mu(\mathcal{S})$, under a volume constraint. Clearly, the optimal $\mathcal{S}$ under this criterion is $r\mathcal{B}$ where =$\mathcal{B}=\{x\in\mathbb{R}^{d}~:~\|x\|\leq r_{\text{eff}}(1)\}$ is a Euclidean ball with radius $r_{\text{eff}}(1)$ chosen such that $\mathrm{vol}(\mathcal{B})=1$, and $r$ is chosen such that $\mathrm{vol}(r\mathcal{B})=r^{d}$ satisfies the required volume constraint. Unfortunately, for $d>1$ the codebook $\mathcal{C}=\Lambda\cap r\mathcal{B}$ loses much of the lattice structure, and does not admit an efficient enumeration, and consequently encoding and decoding require using a lookup table (LUT). QuIP# used this approach with $\Lambda=E_{8}$ (same as we do) and $\mathcal{S}=r\mathcal{B}$. However, this seems to only be possible for quantizing weights and not activations as complexity makes runtime implementation too slow.We note that QuIP# cleverly exploits symmetries of $E_{8}$ to show that an $R=2$ bit quantizer can be implemented using an LUT of size $2^{d\frac{R}{2}}=2^{8}$, but we believe this is still too slow, and furthermore does not naturally extend to different quantization rates.
@@
Voronoi codes/nested lattice codes: In Voronoi codes ConwaySloane83 the same lattice $\Lambda$ is used for both quantization and shaping. In particular, the shaping region is taken as $\mathcal{S}=q\mathcal{V}_{\Lambda}$, where $q=2^{R}$ is an integer. As elaborated below, if $Q_{\Lambda}(x)$ admits an efficient implementation, one can efficiently perform encoding and decoding to the codebook $\mathcal{C}=\Lambda\cap(2^{R}\mathcal{V}_{\Lambda})\cong\Lambda/2^{R}\Lambda$. Moreover, in stark contrast to ball-based shaping, the encoding and decoding complexity does not depend on $R$.
@@
In this work, we use the Gosset lattice ($E_{8}$) that satisfies all these properties. It has a fast decoding algorithm (Algorithm 5), its NSM is $\approx 0.0716821\approx 1.2243\frac{1}{2\pi e}$ agrell2023best , and its Gaussian mass $\mu(r\mathcal{V}_{E_{8}})$ is very close to $\mu(r\mathcal{B})$ (the Gosset lattice has unit covolume, so $\mathrm{vol}(r\mathcal{V}_{E_{8}})=\mathrm{vol}(r\mathcal{B})$). The last point is illustrated in Figure 5, where the large loss for cubic shaping with respect to lattice shaping is also evident. The gap between Voronoi/Ball shaping and cubic shaping becomes much more significant as the dimension increases. This follows since for large $d$ we have that for $X\sim\mu=\mu_{d}$ the $\ell_{\infty}$ norm $\|X\|_{\infty}$ concentrates around $\sqrt{2\ln{d}}$. Thus, for $r<2\sqrt{2\ln{d}}$ we have that $\mu(r\mathrm{CUBE})\to 0$, whereas for any $r>\frac{\sqrt{d}}{r_{\text{eff}}(1)}=\sqrt{2\pi e}(1+o(1))$ we have that $\mu(r\mathcal{B})\to 1$. Note that SpinQuant uses high-dimensional cubic shaping, and therefore its MSE distortion suffers a $O(\ln d)$ multiplicative gap with respect to the optimal distortion.